e.g. 0.1111... = 1/9
So then how would I convert 0.0238095238095238095... (238095 is a recurring decimal)to a fraction?
Thank you!

Answer by stanbon(75887)

convert 0.0238095238095238095... (238095 is a recurring decimal)to a fraction?
----
Let x = 0.0238095238095238095...
Then 10x = 0.238095238095...
And 10^7x = 238095.238095238095
------------------------------
Subtract the 2nd line from the 3rd to get:
(10^7-10)x = 238095
---
x = 238095/[10000000-10] = 238095/9,999,990
=============================================
Cheers,
Stan H.